DEtails.

10:00

Power on

ocam

ok

No light in the CCD47.

Check the AGW. The AGW was stuck.

Call John for help and he reseted the controller and move away the AGw.

pupil wobble

Y has a shift of about 5 pixels in full rotation.

shift y->-1 :'20240426_180636'

rtestart adsec.

pupil shift y-.>-5, offwetY -3

20240426_182544

save tn:20240426_200609.

- the axis still reachong the -100 with an error <-0.4, maybe could work with that error but the looking at the postions of the pupils are too far up.

second TN: 20240426_201309

01:25

meeting with Alessandro and Ricard (Mountain), discussion about to check if the ocam is loose in one direction and if he can check this we can try to clamp the ocam.

Lghts on.

setup for nternal pupils. ->Internal2 setup.

confirmed that the ocam is loose.

Ricardo could move the ocam in x and y direction with a small force, move up/down move pupils horizonatal

Power off ocam and buc2k so Ricardo can take photos with flash.

02:10

power off

Pupil wobble with a shift of Y -5. There are reginos where still saturated with an error ~0.4 pixels. The movement of the wobble is not contibuous with some glitchs.

Before the shift.

Screen Shot 2024-04-26 at 5.18.39 PM.png

Screen Shot 2024-04-26 at 5.19.35 PM.png

-- JuanCarlosGuerra - 26 Apr 2024
Topic revision: r3 - 27 Apr 2024, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback